Find all the values of (1 + 2i)}.
Answer #1

2 = (1+20) 3 bath uigde cube 3. z tai S. 2.Z 2 121 = 2+ y² Here 421, y = 2 So 2²= Jitu 5 so Z3 = 2².2 = 5. (1+2013